# Exclusive quarkonium photoproduction: predictions with the Balitsky-Kovchegov equation including the full impact-parameter dependence

### v1

@@ II.2 The BK equation
Figure 1: The $J/\psi$ production data used for extraction of the model parameters. Measurements by H1 [49] were used for the differential cross section (left panel) and a combination of measurements by ZEUS, H1, and ALICE [50, 51, 49, 52] for the total cross section (right panel).
The BK equation in target rapidity, introduced in Ref. [45], is given by
@@
where we have introduced the constant $\kappa$ that will be used to explore the contribution of the non-linear terms by setting it to either one (the standard BK equation) or zero (the linear part of the equation) [53]. The rapidities $\eta_{j}=\eta-\max\{0,\ln(r^{2}/r^{2}_{j})\}$ introduce a non-local behaviour to the equation and require a prescription to handle the rapidities earlier than the initial rapidity. Here we chose to suppress completely these contributions, that is $N(\eta<0)=0$. See Ref. [42] for a discussion of this issue as well as for the definition of the kernel $K(r,r_{1},r_{2})$. The initial condition is
@@
The values of the parameters are chosen to obtain a reasonable description of the energy evolution of ${\rm J}/\psi$ photoproduction total cross-section and the Mandelstam-$t$ dependence at $W_{\gamma\mathrm{p}}=55$ GeV and $W_{\gamma\mathrm{p}}=78$ GeV as shown in Fig. 1. They are set to: $Q_{s0}^{2}=0.37$ GeV2, $B=3.6$ GeV2, $c=0.5$, and $\gamma=1$. There is also a parameter involved in the running of the strong coupling constant which was set to $C=2.5$ (see Ref. [42]). Parameters were obtained using the following quark masses: $m_{\rm uds}=0.1$ GeV$/c^{2}$, $m_{\rm c}=1.27$ GeV$/c^{2}$ and $m_{\rm b}=4.18$ GeV$/c^{2}$. As an example of the evolution, Fig. 2 shows the dipole amplitude at the initial condition and after evolution to rapidity 10, for the case of parallel dipole-size and impact-parameter vectors. The shape remains similar throughout the evolution but the wavefronts move and the profile in impact parameter grows. Other values of $\theta$ provide a similar picture as discussed in Ref. [42].
@@ II.3 Wave function
As it is customary, we fixed the parameters using the normalisation of the wave functions, their partial width to electron-positron pairs, and the orthogonality of the different states [62, 63]. The parameter values obtained by this procedure using data from PDG 2024 [64] are reported in Table 1. Figure 3 shows the scalar part of the charmonium and bottomonium wave functions for the case of transverse polarisation. Note, that the case of longitudinal polarisation is almost identical to the transverse case so we do not present it here. Although the model of Gaussian distribution of the vector meson in the rest frame boosted to the proper light-cone frame is widely used and accepted  [60, 54, 55, 61, 29, 63], there is a more sophisticated approach for modelling the distribution by solving the Schrödinger equation of $q\bar{q}$ pair with realistic potential in the rest frame of the pair [57, 56]. However, the difference between the quadratic potential (boosted Gaussian model) and more sophisticated potentials is not very big and it can be neglected in the context of these studies.
@@ III.2 Contribution of the non-linear terms
To study the contribution of the non-linear terms to the cross-sections for exclusive diffractive photoproduction of quarkonium states we repeat the previous computations where we set $\kappa=0$ in Eq. (5) and compare with the case $\kappa=1$. Figure 5, left panel, shows the ratio of these two cases for the production of ${\rm J}/\psi$ vector mesons. At the initial condition, there has been no evolution, thus the cross-sections for $\kappa=0$ and $\kappa=1$ are the same and the ratio is very close to one. (There is a small difference due to the effect of the corrections that involve the derivative of the amplitude, see Eq. (4), which is different for the $\kappa=0$ and $\kappa=1$ cases.) As the energy increases and the evolution progresses the ratio increases signalling that the non-linear terms, that enter with a negative sign (see Eq. (5), slow down the evolution more and more. This pattern has a strong dependence on Mandelstam-$t$. This can be intuitively understood from Eq. (2), which shows that the momentum transferred is Fourier conjugate to the impact parameter, so large values of Mandelstam-$t$ sample mainly the region of small impact parameters, where the target is denser. A similar behaviour is observed for all the other vector mesons. To explore this effect a bit further, the right panel of Fig. 5 shows the energy dependence of the ratio of cross sections at large and small values of Mandelstam-$t$. There is a strong dependence of the ratio with energy, with a similar behaviour shown for all vector mesons. The effect of the non-linear corrections increases with energy but seems to be small with respect to the expected experimental uncertainties.

### v2

@@ II.2 The BK equation
Figure 1: The $J/\psi$ production data used for extraction of the model parameters. Measurements by H1 [49] were used for the differential cross section (left panel) and a combination of measurements by ZEUS, H1, ALICE, and LHCb [50, 51, 49, 52, 9] for the total cross section (right panel).
The BK equation in target rapidity was introduced in Ref. [45]. It traces the evolution with the rapidity of the dipole-target amplitude $N(\vec{r},\vec{b};\eta)$ when the parent dipole described by the size and impact parameter vectors $\vec{r}$ and $\vec{b}$, respectively, split into two daughter dipoles characterised by $\vec{r}_{1,2}$ and $\vec{b}_{1,2}$; the equation also includes the case when two dipoles recombine into one. The equation is given by
@@
where we have introduced the constant $\kappa$ that will be used to explore the contribution of the non-linear terms by setting it to either one (the standard BK equation) or zero (the linear part of the equation) [53]. The collinearly improved kernel is
$K(r,r_{1},r_{2})=\frac{\bar{\alpha}_{s}}{2\pi}\frac{r^{2}}{r_{1}^{2}r_{2}^{2}}\bigg{[}\frac{r^{2}}{\min\{r_{1}^{2},r_{2}^{2}\}}\bigg{]}^{\pm\bar{\alpha}_{s}A_{1}},$ (6)
where the first part corresponds to the standard leading log approximation including the running of the strong coupling. The second, exponentiated, part comes from resumming the DGLAP-like single transverse logarithmic contributions [54]. The constant $A_{1}=\frac{11}{12}$ and $J_{1}$ is the Bessel function. Here, $\bar{\alpha}_{s}=\frac{N_{\rm C}}{\pi}\alpha_{s}$ with $\alpha_{s}=\alpha_{s}(\min\{r,r_{1},r_{2}\})$ being the running strong coupling constant evaluated in the scheme with variable number of flavours. The rapidities $\eta_{j}=\eta-\max\{0,\ln(r^{2}/r^{2}_{j})\}$ introduce a non-local behaviour to the equation and require a prescription to handle the rapidities earlier than the initial rapidity. Here we chose to suppress completely these contributions, that is $N(\eta<0)=0$. See Ref. [42] for a discussion of this issue as well as for the definition of the kernel $K(r,r_{1},r_{2})$. The initial condition is
@@
The values of the parameters are chosen to obtain a reasonable description of the energy evolution of ${\rm J}/\psi$ photoproduction total cross-section and the Mandelstam-$t$ dependence at $W_{\gamma\mathrm{p}}=55$ GeV and $W_{\gamma\mathrm{p}}=78$ GeV as shown in Fig. 1. They are set to: $Q_{s0}^{2}=0.37$ GeV2, $B=3.6$ GeV2, $c=0.5$, and $\gamma=1$. There is also a parameter involved in the running of the strong coupling constant which was set to $C=2.5$. These values were updated since our previous work [42], as we extended the dataset for their extraction from solely HERA data to a combination of HERA and LHC, covering a wider energy region. Furthermore, new parameters were obtained for the vector-meson wave function using the latest PDG data, which also needed to be accounted for. The presented BK parameters were obtained using the following quark masses: $m_{\rm uds}=0.1$ GeV$/c^{2}$, $m_{\rm c}=1.27$ GeV$/c^{2}$ and $m_{\rm b}=4.18$ GeV$/c^{2}$. As an example of the evolution, Fig. 2 shows the dipole amplitude at the initial condition and after evolution to rapidity 10, for the case of parallel dipole-size and impact-parameter vectors. The shape remains similar throughout the evolution but the wavefronts move and the profile in impact parameter grows. Other values of $\theta$ provide a similar picture as discussed in Ref. [42].
@@ II.3 Wave function
As it is customary, we fixed the parameters using the normalisation of the wave functions, their partial width to electron-positron pairs, and the orthogonality of the different states [63, 64].
@@
The parameter values obtained by this procedure using data from PDG 2024 [65] are reported in Table 1. Figure 3 shows the scalar part of the charmonium and bottomonium wave functions for the case of transverse polarisation. Note, that the case of longitudinal polarisation is almost identical to the transverse case so we do not present it here. Although the model of Gaussian distribution of the vector meson in the rest frame boosted to the proper light-cone frame is widely used and accepted  [61, 55, 56, 62, 29, 64], there is a more sophisticated approach for modelling the distribution by solving the Schrödinger equation of $q\bar{q}$ pair with realistic potential in the rest frame of the pair [58, 57]. However, the difference between the quadratic potential (boosted Gaussian model) and more sophisticated potentials is not very big and it can be neglected in the context of these studies.
@@ III.2 Contribution of the non-linear terms
To study the contribution of the non-linear terms to the cross-sections for exclusive diffractive photoproduction of quarkonium states we repeat the previous computations where we set $\kappa=0$ in Eq. (5) and compare with the case $\kappa=1$. Figure 5, left panel, shows the ratio of these two cases for the production of ${\rm J}/\psi$ vector mesons. At the initial condition, there has been no evolution, thus the cross-sections for $\kappa=0$ and $\kappa=1$ are the same and the ratio is very close to one. (There is a small difference due to the effect of the corrections that involve the derivative of the amplitude, see Eq. (4), which is different for the $\kappa=0$ and $\kappa=1$ cases.) As the energy increases and the evolution progresses the ratio increases signalling that the non-linear terms, that enter with a negative sign (see Eq. (5)), slow down the evolution more and more. This pattern has a strong dependence on Mandelstam-$t$. This can be intuitively understood from Eq. (2), which shows that the momentum transferred is Fourier conjugate to the impact parameter, so large values of Mandelstam-$t$ sample mainly the region of small impact parameters, where the target is denser. A similar behaviour is observed for all the other vector mesons. To explore this effect a bit further, the right panel of Fig. 5 shows the energy dependence of the ratio of cross sections at large and small values of Mandelstam-$t$. There is a strong dependence of the ratio with energy, with a similar behaviour shown for all vector mesons. The effect of the non-linear corrections increases with energy but seems to be small with respect to the expected experimental uncertainties.